OpenStudy (dessyj1):

Arc Length FInd the arc length within an interval of an equation modeled by y=(ln(csc(4x)))/4

OpenStudy (dessyj1):

\[\ln \left( \csc (4x) \right)\frac{ 1 }{ 4 }\]

OpenStudy (dessyj1):